# On the initial spin period distribution of neutron stars

### v1

@@ Abstract
We derive the initial spin period distribution of neutron stars by studying the population of young pulsars associated with supernova remnants. Our hierarchical Bayesian approach accounts for the measurement uncertainties of individual observations and selection effects. Without correcting for selection effects, as done in previous studies, we find that pulsar initial spin periods follow a Weibull distribution, peaking at 40 ms, which is favoured against the lognormal distribution with a Bayes factor of 200. The known selection effects in radio pulsar surveys, including pulse broadening and a period-dependent beaming fraction, have been quantitatively investigated. We show that, based on measurements of pulsar luminosity and spin period from the ATNF Pulsar Catalogue, the impact of pulse broadening on the inference of pulsar initial period distribution is likely to be insignificant. Correcting for the beaming selection effect, a Weibull distribution remains to be the preferred model, while its peak slightly shifts to longer periods at 50 ms. Our method will prove useful in constraining the birth properties of neutron stars in the Square Kilometre Array era.
@@
software: BILBY (Ashton et al., 2019), DYNESTY (Speagle, 2020), ChainConsumer (https://github.com/Samreay/ChainConsumer.git)
@@ 1 Introduction
Pulsars are fast rotating, highly magnetized neutron stars (NSs). Since the first discovery (Hewish et al., 1968), the number of observed pulsars has grown to over 3500 (see the Australia Telescope National Facility (ATNF) Pulsar Catalogue [^1] https://www.atnf.csiro.au/research/pulsar/psrcat/, Manchester et al., 2005), a majority of which are detected in the radio band. Among them, pulsars observed in association with supernova remnants (SNRs) are of particular interest, since their ages are independently informed by observations of SNRs. The Crab pulsar is the best known example. Firmly established as the remnant star of supernova 1054, its initial spin period is around 20 ms, close to its current spin period of 33 ms. This is widely used as a proxy for pulsar initial spin periods in the community (e.g., Johnston & Karastergiou, 2017).
@@
The astrophysical processes that give rise to NS spins are poorly understood. A range of spin periods from milliseconds to seconds are predicted in a variety of processes during supernova explosions. Newborn NSs could inherit the angular momentum of progenitor stars from the collapsing iron cores (Heger et al., 2005; Ott et al., 2006), where the angular momentum transport and mass loss in single stars play a significant role (Fuller et al., 2014, 2015, 2019; Ma & Fuller, 2019; Eggenberger et al., 2019). For instance, Ott et al. (2006) found that an NS could be born with periods of tens to hundreds of milliseconds if the spin periods of iron cores are around 50--100 s. NS spins could also stem from the natal kicks caused by the asymmetric mass ejection and anisotropic neutrino emission (e.g., Spruit & Phinney 1998; Ng & Romani 2007; Janka et al. 2022; Coleman & Burrows 2022; Fragione & Loeb 2023). In particular, Spruit & Phinney (1998) showed that an off-center kick at a speed of several hundreds km///s could lead to a spin period as short as tens of milliseconds. Other processes that might play a role in producing NS spins include hydrodynamic instabilities during the supernova explosions such as the standing accretion shock instability (e.g., Blondin & Mezzacappa 2007; Guilet & Fernández 2014; Kazeroni et al. 2016), the anisotropic accretion of angular momentum during the pre-explosion phase (Wongwathanarat et al., 2013) and post-explosion phase (e.g., Janka et al., 2022). Therefore, the initial period distribution of NSs is a powerful probe into astrophysical processes during their formation.
@@
Thus far, around a hundred pulsars have been detected in association with SNRs. They are considered to be young (with a typical age less than $\sim 10^{5}$ years), and therefore provide useful insights into the birth properties of NSs, such as the spin period, magnetic field, spatial velocity, and inclination angle between the spin and magnetic axes (Popov & Turolla, 2012; Malov, 2021; Igoshev et al., 2022). To properly derive the initial distribution of pulsar parameters with the observed pulsar-SNR population, two approaches can be used. First, by performing population syntheses (Emmering & Chevalier, 1989; Faucher-Giguère & Kaspi, 2006; Bates et al., 2014; Gullón et al., 2014; Cieślar et al., 2020; Dirson et al., 2022), one develops a model pulsar population based on plausible assumptions about their initial properties and time evolution, and then runs the synthetic population through some mock pulsar surveys with an aim to reproduce the observed pulsar sample. The model that best reproduces observations is considered to be a good model. In the second approach, one extrapolates, for example, the initial periods of pulsars back from their measured spins and age estimates with assumptions about spin evolution. Then, statistical inference is performed to fit the distribution of initial periods by incorporating selection effects and measurement uncertainties in individual observations. The major challenge in this approach is an appropriate quantification of the possible selection effects that would bias the observed distribution from the true population (e.g., Lorimer et al. 2006, 2015). Igoshev et al. (2022) performed maximum-likelihood estimates for the distributions of initial periods and magnetic fields of NSs using a sample of Galactic pulsars detected in association with SNRs. However, the selection effects were only qualitatively discussed.
@@ 2.1 Pulsars in supernova remnants
Figure 1: Upper panel: The $P$-$\dot{P}$ diagram of the observed pulsars. The stars denote the pulsars detected within SNRs, with CCOs and magnetars highlighted in yellow and cyan, respectively. The age and magnetic field strength ($B_{\rm dip}$) are calculated in the standard magneto-dipole model. Note that three pulsar-SNR pairs detected in the Large Magellanic Cloud are also included. Lower panel: The distributions of the observed spin periods and initial periods of pulsars associated with SNRs. The green line denotes the posterior predictive distribution of initial period derived from the Weibull distribution (see section 3.2.1), with the shaded areas being the 90% credible interval. The gray dashed line is the best-fitting result obtained by Igoshev et al. (2022) assuming a log-normal distribution.
@@
We show the pulsar $P$-$\dot{P}$ diagram in the upper panel of Figure 1. The NSs associated with SNRs are denoted by stars: magnetars (cyan) populate the slowest and strongest magnetic filed regions among all species of neutron stars; normal pulsars and two CCOs are highlighted in red and yellow, respectively. In the lower panel, we show the observed spin period distribution of the pulsars included in our sample.
@@ 2.2 Spin-down evolution of pulsars
To avoid imaginary or zero periods ($P^{2}_{0}\leq 0$) when taking $n=3$ in Equation (3), $\tau_{\rm snr}<\tau_{\rm c}$ must be satisfied, which is an additional selection criterion for our sample [^4] There are ten (out of 45) pulsars fall into this condition in the sample used by Igoshev et al. (2022), in which case $P_{0}=2$ ms is set.. There are six pulsars with $n$ measurements in this sample. In the lower panel of Figure 1, we show the derived initial periods of the selected pulsars (blue histogram).
@@ 3.1 Method
Figure 2: The posterior predictive distributions of NS initial spin periods. Upper panel: The Weilbull distribution derived without correcting for selection effects and by adopting fixed (blue) and random (orange) values of braking index. Lower panel: The Weilbull distribution derived by correcting the beaming fraction using the LM88 (green) and TM98 (gray) models (see Section 3.2.3 for details). Shaded areas indicate the $90\%$ credible intervals.
@@ 3.2.1 Without correcting selection effects
Figure 2 plots the posterior predictive distribution (PPD) of initial spin period for the Weibull model:
@@
The shaded areas denotes the 90% credible interval. The Weibull distribution indicates that the majority of the observed pulsars were initially fast spinning, with initial periods peaking at $\sim$40 ms. The 90% credibility upper limit is constrained to be $P_{0}^{90\%}\approx 0.5$ s. In the lower panel of Figure 1, we compare the initial period distribution inferred from the Weibull distribution (in green) with the best-fit result (in gray) obtained by Igoshev et al. (2022). We obtain a narrower initial period distribution. The difference of the results can be attributed to different data selection criteria, especially the treatment of the observed pulsars with imaginary initial periods (see Section 2.2), and different analysis method.
@@ 3.2.3 Correcting for selection effects
Figure 4: Panel (a): The sensitivity of Parkes Multibeam Survey. The minimum detectable flux density ($S_{\rm min}$) is computed using Equation (C6). Panel (b): The detectability of radio pulsars. The solid and dashed lines account for the beaming fraction of the LM88 and TM98 models, respectively, see Appendix D. The gray dots denote the detectability obtained with the inverse of scale factors (see Appendix C).
@@
The intrinsic radio-luminosity function remains poorly constrained (see Posselt et al. 2023 and references therein). In Figure 3, we show that there is no significant dependence of the observed radio (pseudo) luminosity on the period for radio pulsars with periods $P>10$ ms and $B_{\rm dip}\geq 10^{11}$ G in the ATNF Catalogue. We adopt the observed luminosity distribution to compute the scale factors [see Equation (C4)]. We calculate the scale factors as a function of $P$ based on the sensitivity of the Parkes Multibeam Survey66615 out of 39 in our sample were detected in this survey. The remaining pulsars come from the Fermi Gamma-ray Observatory blind survey, Green Bank Telescope, Arecibo surveys, and other searches.. The sensitivity curve of this survey is shown in panel (a) of Figure 4. (The calculation details are described in Appendix C.) Shorter-period and farther pulsars are more susceptible to pulse smearing in the interstellar medium, resulting in lower sensitivities in their detection. However, the sensitivity becomes almost constant for pulsars with longer periods (e.g., $P\gtrsim 0.1$ s). The detection probability, estimated from the inverse of the scale factor, is represented by gray dots in the panel (b). The detectability grows slightly with periods for $P\lesssim 0.1$ s, which corresponds to the effect of the sensitivity limit on radio observations shown in panel (a). Overall, the influence of pulse smearing on the shape of the initial period distribution are small, which is consistent with the fact that pulsar luminosity does not depend on spin periods as shown in Figure 3.
Pulsars with shorter periods are found to possess wider radio beams, resulting in higher detectability in pulsar surveys (Lyne & Manchester, 1988; Tauris & Manchester, 1998). Here, we consider two empirical models, LM88 (Lyne & Manchester, 1988) and TM98 (Tauris & Manchester, 1998), to quantify the period-beaming fraction correlation (see Appendix D for details), which is plotted in the panel (b) of Figure 4. Incorporating beaming fraction corrections, we obtain similar constraints of population parameters for both models adopted, see Table 1. The Bayes factors given in Table 2 indicate that the Weibull distribution is still the best description of pulsar initial period distribution. In the lower panel of Figure 2, one can see a slight shift towards longer periods, with a peak at around 50 ms. The 90% credibility upper limit is constrained to be $P_{0}^{90\%}\approx 0.7$ s (0.8 s) for the LM88 (TM98) model.
@@ 4 Conclusions
Z.-H. Zhu is supported by the National Natural Science Foundation of China under Grants Nos. 12021003, 11920101003 and 11633001, and the Strategic Priority Research Program of the Chinese Academy of Sciences, Grant No. XDB23000000. ZQY is supported by the National Natural Science Foundation of China under Grant No. 12305059. XJZ is supported by the National Natural Science Foundation of China (Grant No. 12203004). ZCC is supported by the National Natural Science Foundation of China (Grant No. 12247176 and No. 12247112) and the China Postdoctoral Science Foundation Fellowship No. 2022M710429.
@@ Appendix A Catalogue of neutron stars associated with SNRs
45 ${}^{\spadesuit}$ G205.5+0.5 30--150 59.2 0.297 7.96E--14 4.92E+12 - N
46 ${}^{\spadesuit}$ Monogem Ring 86--170 111 0.385 5.49E--14 4.65E+12 - N
@@
57 ${}^{\dagger}$ G333.9+00.0 0--6 24.7 4.327 2.78E--12 1.11E+14 - N
58 ${}^{\spadesuit}$ G336.1--0.2 - 19.9 0.814 6.49E--13 2.33E+13 - N
59 ${}^{\dagger}$ G337.0--00.1 5--5 - 2.595 - - - N
60 J1640--4631 G338.3--00.0 1--8 3.35 0.206 9.76E--13 1.44E+13 3.15$\pm$0.03${}^{(7)}$ N
61 J1640--4631 G338.5--00.1 1.1--17.0 3.35 0.206 9.76E--13 1.44E+13 3.15$\pm$0.03${}^{(7)}$ N
62 ${}^{\spadesuit}$ G341.2+0.9 - 32.5 0.232 1.13E--13 5.17E+12 - N
63 ${}^{\spadesuit}$ G343.1--2.3 - 17.5 0.102 9.30E--14 3.12E+12 - N
64 ${}^{\dagger}$ G348.7+00.3 0.65--16.8 1.03 3.825 5.88E--11 4.80E+14 - N
65 ${}^{\spadesuit}$ G352.2--0.1 - 14.5 1.11 1.22E--12 3.72E+13 - N
66 ${}^{\spadesuit}$ G343.0--6.0 - 80.4 0.83 1.64E--13 1.18E+13 - N
67 ${}^{\dagger}$${}^{\spadesuit}$ G10.0--0.3 - 0.218 7.556 5.49E--10 2.06E+15 - N
@@
86 ${}^{\dagger}$ G057.2+00.8 16--95 - 3.245 - - - N
87 ${}^{\spadesuit}$ G76.9+1.0 - 8.94 0.049 8.61E--14 2.07E+12 - N
88 ${}^{\dagger}$ G109.1--01.0 8.8--14 235 6.979 4.71E--13 5.80E+13 - N
footnotetext: The objects known as magnetars in the McGill Online Magnetar Catalog.footnotetext: The objects known as central compact objects.footnotetext: The objects are not included in the sample of Igoshev et al. (2022).

### v2

@@ Abstract
We derive the initial spin period distribution of neutron stars by studying the population of young pulsars associated with supernova remnants. Our hierarchical Bayesian approach accounts for the measurement uncertainties of individual observations and selection effects. Without correcting for selection effects, as done in previous studies, we find that pulsar initial spin periods follow a Weibull distribution, peaking at 40 ms, which is favoured against the lognormal distribution with a Bayes factor of 200. The known selection effects in radio pulsar surveys, including pulse broadening and a period-dependent beaming fraction, have been quantitatively investigated. We show that, based on measurements of pulsar luminosity and spin period from the ATNF Pulsar Catalogue, the impact of pulse broadening on the inference of pulsar period distribution is likely to be insignificant. Correcting for the beaming selection effect, a Weibull distribution remains to be the preferred model, while its peak slightly shifts to longer periods at 50 ms. Our method will prove useful in constraining the birth properties of neutron stars in the Square Kilometre Array era.
@@
software: BILBY (Ashton et al., 2019), DYNESTY (Speagle, 2020), ChainConsumer (Hinton, 2016; Hinton et al., 2024), Scipy (Virtanen et al., 2021).
@@ 1 Introduction
Pulsars are fast rotating, highly magnetized neutron stars (NSs). Since the first discovery (Hewish et al., 1968), the number of observed pulsars has grown to over 3500 (see the ATNF Pulsar Catalogue [^1] https://www.atnf.csiro.au/research/pulsar/psrcat/, Manchester et al., 2005), a majority of which are detected in the radio band. Among them, pulsars observed in association with supernova remnants (SNRs) are of particular interest, since their ages are independently informed by observations of SNRs. The Crab pulsar is the best known example. Firmly established as the remnant star of supernova 1054, its initial spin period is around 20 ms, close to its current spin period of 33 ms. This is widely used as a proxy for pulsar initial spin periods in the community (e.g., Johnston & Karastergiou, 2017).
@@
The astrophysical processes that give rise to NS spins are poorly understood. A range of spin periods from milliseconds to seconds are predicted in a variety of processes during supernova explosions. Newborn NSs could inherit the angular momentum of progenitor stars from the collapsing iron cores (Heger et al., 2005; Ott et al., 2006), where the angular momentum transport and mass loss in single stars play a significant role (Fuller et al., 2014, 2015, 2019; Ma & Fuller, 2019; Eggenberger et al., 2019; Hu et al., 2023). For instance, Ott et al. (2006) found that an NS could be born with periods of tens to hundreds of milliseconds if the spin periods of iron cores are around 50--100 s. NS spins could also stem from the natal kicks caused by the asymmetric mass ejection and anisotropic neutrino emission (e.g., Spruit & Phinney 1998; Ng & Romani 2007; Janka et al. 2022; Coleman & Burrows 2022; Fragione & Loeb 2023; Burrows et al. 2023). In particular, Spruit & Phinney (1998) showed that an off-center kick at a speed of several hundreds km///s could lead to a spin period as short as tens of milliseconds. Other processes that might play a role in producing NS spins include hydrodynamic instabilities during the supernova explosions such as the standing accretion shock instability (e.g., Blondin & Mezzacappa 2007; Guilet & Fernández 2014; Kazeroni et al. 2016), the anisotropic accretion of angular momentum during the pre-explosion phase (Wongwathanarat et al., 2013) and post-explosion phase (e.g., Janka et al., 2022). Therefore, the initial period distribution of NSs is a powerful probe into astrophysical processes during their formation.
@@
Thus far, around a hundred pulsars have been detected in association with SNRs. They are considered to be young (with a typical age less than $\sim 10^{5}$ years), and therefore provide useful insights into the birth properties of NSs, such as the spin period, magnetic field, spatial velocity, and inclination angle between the spin and magnetic axes (Popov & Turolla, 2012; Malov, 2021; Igoshev et al., 2022). To properly derive the initial distribution of pulsar parameters with the observed pulsar-SNR population, two approaches can be used. First, by performing population syntheses (Emmering & Chevalier, 1989; Faucher-Giguère & Kaspi, 2006; Bates et al., 2014; Gullón et al., 2014; Cieślar et al., 2020; Dirson et al., 2022), one develops a model pulsar population based on plausible assumptions about their initial properties and time evolution, and then runs the synthetic population through some mock pulsar surveys with an aim to reproduce the observed pulsar sample. The model that best reproduces observations is considered to be a good model. In the second approach, one extrapolates, for example, the initial periods of pulsars back from their measured spins and age estimates with assumptions about spin evolution (e.g., Xu et al. 2023). Then, statistical inference is performed to fit the distribution of initial periods by incorporating selection effects and measurement uncertainties in individual observations. The major challenge in this approach is an appropriate quantification of the possible selection effects that would bias the observed distribution from the true population (e.g., Lorimer et al. 2006, 2015). Igoshev et al. (2022) performed maximum-likelihood estimates for the distributions of initial periods and magnetic fields of NSs using a sample of Galactic pulsars detected in association with SNRs. However, the selection effects were only qualitatively discussed.
@@ 2.1 Pulsars in supernova remnants
Figure 1: The $P$-$\dot{P}$ diagram of the observed pulsars. The stars denote the pulsars detected within SNRs, with CCOs and magnetars highlighted in yellow and cyan, respectively. The age and magnetic field strength ($B_{\rm dip}$) are calculated in the standard magneto-dipole model. The data set of this work is available at https://github.com/Shen-Shi/NS_Birth_Population.
@@
We show the pulsar $P$-$\dot{P}$ diagram in Figure 1. The NSs associated with SNRs are denoted by stars: magnetars (cyan) populate the slowest and strongest magnetic filed regions among all species of neutron stars; normal pulsars and two CCOs are highlighted in red and yellow, respectively. In the upper panel of Figure 2, we show the observed spin period distribution of the pulsars (orange histogram) included in our sample.
Figure 2: Upper panel: The distributions of the observed spin periods (orange) and initial periods (blue) of pulsars associated with SNRs. The green (gray) line denotes the posterior predictive distribution of initial period derived from the Weibull (log-normal) distribution (see section 3.2.1), with the shaded area indicating the 90% credible interval. Lower panel: The cumulative distributions of the initial periods and the Weibull and log-normal model are plotted with the same colors as the upper panel.
@@ 2.2 Spin-down evolution of pulsars
To avoid imaginary or zero periods ($P^{2}_{0}\leq 0$) when taking $n=3$ in Equation (3), $\tau_{\rm snr}<\tau_{\rm c}$ must be satisfied, which is an additional selection criterion for our sample [^4] There are nine (out of 45) pulsars fall into this condition in the sample used by Igoshev et al. (2022), in which case $P_{0}=2$ ms is set.. There are six pulsars with $n$ measurements in this sample. In the upper panel of Figure 2, we show the derived initial periods of the selected pulsars (blue histogram).
@@ 3.1 Method
Figure 3: The posterior predictive distributions of NS initial spin periods. Upper panel: The Weilbull distribution derived without correcting for selection effects and by adopting fixed (blue) and random (orange) values of braking index. Lower panel: The Weilbull distribution derived by correcting the beaming fraction using the LM88 (green) and TM98 (gray) models (see Section 3.2.3 for details). The shaded areas indicate the $90\%$ credible intervals.
@@ 3.2.1 Without correcting selection effects
Figure 2 plots the posterior predictive distribution (PPD) defined as:
@@
The PPD of the Weibull and LOGN model are shown in green and grey, respectively. The shaded areas denotes the 90% credible interval. The Weibull distribution, which is the preferred model, indicates that the majority of the observed pulsars were initially fast spinning, with initial periods peaking at $\sim$40 ms. The 90% credibility upper limit is constrained to be $P_{0}^{90\%}\approx 0.5$ s. The cumulative distributions of the Weibull and LOGN are shown in the lower panel of Figure 2, where one can see that the Weibull model provides a better fit than the LOGN. In Appendix C, we compare the initial period distribution inferred with the LOGN model with that obtained by Igoshev et al. (2022).
Table 3: The results of two-sample Kolmogorov-Smirnoff test for different distribution functions.
Model Test statistic p-value
Weibull 0.070 0.686
TOPL 0.133 0.053
LOGN 0.115 0.129
GAMMA 0.207 $3.18\times 10^{-4}$
GS 0.437 $<10^{-4}$
To additionally provide a quantitative measure of the choices of the distribution functions, we perform a two-sample Kolmogorov–Smirnov test implemented by the scipy package (Virtanen et al., 2021) [^6] Two-sample Kolmogorov–Smirnov test evaluates whether two samples are subject to the same underlying distribution or not.. The compared samples are (1) the initial periods from 39 pulsars (as presented in section 2) and (2) 100 synthetic initial periods drawn from each distribution with the median values of the parameters in Table 1; see the column “no selection ($n=3$)”. The values of test statistic, defined as the largest deviation between the two compared cumulative distributions, and p-value are presented in Table 3. We observe from Table 3 that the Weibull is the best model; the TOPL and LOGN models come next; and the GAMMA and Gaussian models are disfavoured. A Kolmogorov–Smirnov test repeated with 200 synthetic initial periods gives consistent results. Overall, the Kolmogorov–Smirnov test results support Bayesian model selection results.
@@ 3.2.3 Correcting for selection effects
Figure 5: Panel (a): The sensitivity of Parkes Multibeam Survey. The minimum detectable flux density ($S_{\rm min}$) is computed using Equation (D6). Panel (b): The detectability of radio pulsars. The solid and dashed lines account for the beaming fraction of the LM88 and TM98 models, respectively, see Appendix E. The gray dots denote the detectability obtained with the inverse of scale factors (see Appendix D).
@@
The intrinsic radio-luminosity function remains poorly constrained (see Posselt et al. 2023 and references therein). In Figure 4, we show that there is no significant dependence of the observed radio (pseudo) luminosity on the period for radio pulsars with periods $P>10$ ms and $B_{\rm dip}\geq 10^{11}$ G in the ATNF Catalogue. We adopt the observed luminosity distribution to compute the scale factors [see Equation (D4)]. We calculate the scale factors as a function of $P$ based on the sensitivity of the Parkes Multibeam Survey77715 out of 39 in our sample were detected in this survey. The remaining pulsars come from the Fermi Gamma-ray Observatory blind survey, Green Bank Telescope, Arecibo surveys, and other searches.. The sensitivity curve of this survey is shown in panel (a) of Figure 5. (The calculation details are described in Appendix D.) Shorter-period and farther pulsars are more susceptible to pulse smearing in the interstellar medium, resulting in lower sensitivities in their detection. However, the sensitivity becomes almost constant for pulsars with longer periods (e.g., $P\gtrsim 0.1$ s). The detection probability, estimated from the inverse of the scale factor, is represented by gray dots in the panel (b). The detectability grows slightly with periods for $P\lesssim 0.1$ s, which corresponds to the effect of the sensitivity limit on radio observations shown in panel (a). Overall, the influence of pulse smearing on the shape of the initial period distribution are small, which is consistent with the fact that pulsar luminosity does not depend on spin periods as shown in Figure 4.
Pulsars with shorter periods are found to possess wider radio beams, resulting in higher detectability in pulsar surveys (Lyne & Manchester, 1988; Tauris & Manchester, 1998). Here, we consider two empirical models, LM88 (Lyne & Manchester, 1988) and TM98 (Tauris & Manchester, 1998), to quantify the period-beaming fraction correlation (see Appendix E for details), which is plotted in the panel (b) of Figure 5. Incorporating beaming fraction corrections, we obtain similar constraints of population parameters for both models adopted, see Table 1. The Bayes factors given in Table 2 indicate that the Weibull distribution is still the best description of pulsar initial period distribution. In the lower panel of Figure 3, one can see a slight shift towards longer periods, with a peak at around 50 ms. The 90% credibility upper limit is constrained to be $P_{0}^{90\%}\approx 0.7$ s (0.8 s) for the LM88 (TM98) model.
@@ 4 Conclusions
We thank the anonymous referee for valuable comments on the manuscript. Z.-H. Zhu is supported by the National Natural Science Foundation of China under Grants Nos. 12021003, 11920101003 and 11633001, and the Strategic Priority Research Program of the Chinese Academy of Sciences, Grant No. XDB23000000. ZQY is supported by the National Natural Science Foundation of China under Grant No. 12305059. XJZ is supported by the National Natural Science Foundation of China (Grant No. 12203004) and by the Fundamental Research Funds for the Central Universities. ZCC is supported by the National Natural Science Foundation of China (Grant No. 12247176 and No. 12247112) and the innovative research group of Hunan Province under Grant No. 2024JJ1006.
@@ Appendix A Catalogue of neutron stars associated with SNRs
45 J0633+0632 G205.5+0.5 30--150 59.2 0.297 7.96E--14 4.92E+12 - N
46 J0659+1414 Monogem Ring 86--170 111 0.385 5.49E--14 4.65E+12 - N
@@
57 J1622--4950 G333.9+00.0 0--6 24.7 4.327 2.78E--12 1.11E+14 - N
58 J1632--4818 G336.1--0.2 - 19.9 0.814 6.49E--13 2.33E+13 - N
59 J1635--4735 G337.0--00.1 5--5 - 2.595 - - - N
60 J1640--4631 G338.3--00.0 1--8 3.35 0.206 9.76E--13 1.44E+13 3.15$\pm$0.03(7) N
61 J1640--4631 G338.5--00.1 1.1--17.0 3.35 0.206 9.76E--13 1.44E+13 3.15$\pm$0.03(7) N
62 J1646--4346 G341.2+0.9 - 32.5 0.232 1.13E--13 5.17E+12 - N
63 J1709--4429 G343.1--2.3 - 17.5 0.102 9.30E--14 3.12E+12 - N
64 J1714--3810 G348.7+00.3 0.65--16.8 1.03 3.825 5.88E--11 4.80E+14 - N
65 J1726--3530 G352.2--0.1 - 14.5 1.11 1.22E--12 3.72E+13 - N
66 J1731--4744 G343.0--6.0 - 80.4 0.83 1.64E--13 1.18E+13 - N
67 J1808--2024 G10.0--0.3 - 0.218 7.556 5.49E--10 2.06E+15 - N
@@
86 J1935+2154 G057.2+00.8 16--95 - 3.245 - - - N
87 J2022+3842 G76.9+1.0 - 8.94 0.049 8.61E--14 2.07E+12 - N
88 J2301+5852 G109.1--01.0 8.8--14 235 6.979 4.71E--13 5.80E+13 - N
footnotetext: The objects known as magnetars in the McGill Online Magnetar Catalog.
footnotetext: The objects known as central compact objects.
footnotetext: The objects are not included in the sample of Igoshev et al. (2022).